OpenStudy (anonymous):

A fossilized animal that was killed by a volcanic eruption has 6.25 percent of its original amount of carbon-14 remaining; this radioisotope has a half life of 5730 years. How long ago did the volcano erupt?

OpenStudy (anonymous):

Solution 1 \[6.25percent -> 0.0625 -> 1/16 -> (1/2)^4\] Half life: 5730 years Time = 5730*4 =22920 Solution 2 r: rate of carbon-14 diminishing \[0.0625=r^t\] \[0.5=r^{5730}\]\[r=0.5^{1/5730}\] \[0.0625=(0.5^{1/5730})^t\] \[log(1/16)=tlog(0.5^{1/5730})\] \[t=log(1/16)/log(0.5^{1/5730})\] 22920